### 08 Jan 2024 v1

# Intra-night optical variability of peculiar narrow-line Seyfert 1 galaxies with enigmatic jet behavior

Abstract
Variability studies of active galactic nuclei are a powerful diagnostic tool in understanding the physical processes occurring in disk-jet regions, unresolved by direct imaging with currently available techniques. Here, we report the first attempt to systematically characterize intra-night optical variability (INOV) for a sample of seven apparently radio-quiet narrow-line Seyfert 1 galaxies (RQNLSy1s) that had shown recurring flaring at 37 GHz in the radio observations at Metsähovi Radio Observatory (MRO), indicating the presence of relativistic jets in them, but no evidence for relativistic jets in the recent radio observations of Karl G. Jansky Very Large Array (JVLA) at 1.6, 5.2, and 9.0 GHz. We have conducted a total of 28 intra-night sessions, each lasting 3 hrs for this sample, resulting in an INOV duty cycle (20%) similar to that reported for -ray-NLSy1s (DC 25% - 30%), that display blazar-like INOV. This in turn infers the presence of relativistic jet in our sample sources. Thus, it appears that even lower-mass (M10 M) RQNLSy1 galaxies can maintain blazar-like activities. However, we note that the magnetic reconnection in the magnetosphere of the black hole can also be a viable mechanism to give rise to the INOV from these sources.
keywords:
surveys – galaxies: active – galaxies: jets – -ray-galaxies: photometry – galaxies: Seyfert – gamma-rays: galaxies.1 Introduction
Accretion of gas around the central supermassive black holes (SMBH) of masses is the main powering mechanism of active galactic nuclei (AGNs) that makes them the most energetic objects in the universe with integrated luminosities reaching up to 10 erg s (Koratkar & Blaes, 1999; Bischetti et al., 2017). Among the different observational characteristics, variability on different timescales ranging from minutes to decades across the electromagnetic spectrum is being used as one of the defining characteristics of AGNs (Gaskell & Klimek, 2003; Padovani et al., 2017). Variability studies in AGNs play an important role in understanding the physical processes occurring in these objects. For instance, AGN variability has been used to probe emission mechanisms occurring on physical scales that are unresolved by currently available telescopes/facilities, and also used to investigate the spin and mass of the central SMBH (Urry & Padovani, 1995; Wagner & Witzel, 1995; Ulrich et al., 1997; Zensus, 1997; Cackett et al., 2013; McHardy et al., 2014; Emmanoulopoulos et al., 2014). It is widely believed that the optically thick accretion disk surrounding the SMBH is primarily responsible for the optical emission from AGNs (Shakura & Sunyaev, 1973), but the physical processes producing optical variability are not clearly understood. In a tiny subset of AGNs called blazars, the optical variability is thought to originate due to relativistic boosting of small fluctuations arising through the turbulence of plasma in the jet (e.g., see Marscher & Travis, 1991; Goyal et al., 2012; Calafut & Wiita, 2015). Short-term optical flux variability of AGNs from minutes to hours is known as ‘Intra-Night Optical Variability’ (INOV, Gopal-Krishna et al., 2003).
Among the INOV study of different luminous classes covered by Goyal et al. (2013b), strong INOV with duty cycle (DC) above 30 percent is exhibited by high optical polarization core-dominated quasars, and TeV blazars (both are radio-loud111Radio-loudness is parameterized by the ratio of radio to optical flux densities at 5 GHz and at 4400Å, respectively, with RL 10 and 10 for radio-quiet and radio-loud AGNs, respectively (e.g. see, Kellermann et al., 1989).). This suggests that strong INOV can be an effective tracer of jet activity in AGNs. On the other hand, a low level of INOV DC 10 percent observed in radio-quiet quasars (e.g., see Goyal et al., 2013b) is thought to originate either from its weak jet (e.g., Kellermann et al., 2016) and/or transient shocks or ‘hot spots’ in the accretion disk around the SMBH (Chakrabarti & Wiita, 1993; Mangalam & Wiita, 1993).
Among the different sub-classes of lower-luminosity AGNs, a handful of the radio-loud minority of Narrow-line Seyfert 1 (NLSy1) galaxies, marked by detection in -ray band, exhibit comparable DC ( 30 percent) to the jetted class of AGN (e.g., see Paliya et al., 2013; Ojha et al., 2021). NLSy1s are characterized by smaller width of Balmer emission lines with FWHM(H 2000 km s and flux ratio of [O 3 (Osterbrock & Pogge, 1985; Shuder & Osterbrock, 1981; Goodrich et al., 1989). The ample majority of NLSy1s are radio-quiet, only a tiny fraction 7% is radio-loud and likely to harbor relativistic jets (Komossa et al., 2006; Singh & Chand, 2018). The presence of relativistic jets has been proven in gamma-ray-detected RLNLSy1s (-RLNLSy1s, e.g., Giroletti et al., 2011; Caccianiga et al., 2015; Jarvela et al., 2022). Unlike the general population of RLNLSy1s, -RLNLSy1s show much stronger INOV similar to blazars, suggesting a connection of relativistic jets with their strong INOV. Despite many similar characteristics such as flat radio spectrum, high brightness temperature, superluminal motion, and -ray detection (Yuan et al., 2008; Abdo et al., 2009a, b, c; Foschini et al., 2010; Foschini, 2011; D’Ammando et al., 2012; Berton et al., 2017; D’Ammando et al., 2015; Yao et al., 2015; Lister, 2018; Paliya et al., 2018; Yang et al., 2018; Yao et al., 2019; Foschini et al., 2021, 2022; Li et al., 2023) to blazar class of AGN, an important difference between NLSy1s and blazars is an order of less massive black hole for NLSy1s (, Grupe & Mathur, 2004; Deo et al., 2006; Peterson, 2011). This makes them to harbor less powerful jets (Angelakis et al., 2015; Gu et al., 2015; Fuhrmann et al., 2016; Paliya, 2019).
Recently, a sample of seven radio-quiet and/or radio-silent (never detected in radio) NLSy1s exhibited recurring flaring at 37 GHz in the radio observations at MRO (see Lähteenmäki et al., 2018). However, when these RQNLSy1s were observed with JVLA in A configuration at three different frequencies, 1.6 GHz, 5.2 GHz, and 9.0 GHz, no hints of relativistic jet was detected from them (see Berton et al., 2020b). Taking advantage of INOV, here, we investigate the presence of relativistic jets in these NLSy1s by performing intranight optical monitoring of these seven RQNLSy1s with 12.5m ground-based optical telescopes.
The format of this paper is as follows. In Sect. 2, we describe our optical monitoring and data reduction procedure. Sect. 3 provides details of statistical analysis methods. The main results of this work followed by discussion are given in Sect. 4.
S DSS name | R-mag | |||
M | ||||
J102906.69555625.2 | 19.10 | 0.45 | – | 7.33 |
J122844.81501751.2 | 17.80 | 0.26 | – | 6.84 |
J123220.11495721.8 | 16.90 | 0.26 | – | 7.30 |
J150916.18613716.7 | 18.60 | 0.20 | – | 6.66 |
J151020.06554722.0 | 17.80 | 0.15 | – | 6.67 |
J152205.41393441.3 | 13.10 | 0.08 | 02 | 5.97 |
J164100.10345452.7 | 16.00 | 0.16 | 13 | 7.15 |
SDSS name of RQNLSy1s. | ||||
R-band magnitude of NLSy1s taken from Monet (1998). | ||||
Redshift of the RQNLSy1s taken from Lähteenmäki et al. (2018) . | ||||
Both radio-loudness or radio-silent | ||||
and black hole masses of current sources are taken from | ||||
Järvelä et al. (2015) and Lähteenmäki et al. (2018). In both | ||||
articles, black hole masses were estimated following the | ||||
FWHM(H) - luminosity mass scaling relation, given | ||||
by Greene & Ho (2005). |
Telescope (s) | No. of sessions | Detector (s) | Readout speed | Field of view | Readout | Gain | Focal ratio | Pixel size | Plate scale |
(arcmin) | noise | (e | of | of CCD | of CCD | ||||
(e) | /ADU) | telescope | (m) | ( /pixel ) | |||||
1.04m ST | 05 | 4k4k | 100 kHz | 15.7015.70 | 3.0 | 10.0 | f/13 | 15.0 | 0.23 |
1.30m DFOT | 10 | 2k2k | 100 kHz | 18.2718.27 | 7.5 | 2.0 | f/4 | 13.5 | 0.53 |
1.20m Mt Abu | 12 | 1k1k | 50 kHz | 5.215.21 | 5.0 | 5.0 | f/13 | 13.0 | 0.30 |
2.50m Mt Abu | 01 | 4k4k | 100 kHz | 10.0010.00 | 2.1 | 3.0 | f/8 | 15.0 | 0.15 |
Sampurnand Telescope (ST), Devasthal Fast Optical Telescope (DFOT), Mount Abu telescope. |
2 Optical intra-night monitoring and data reduction
Optical telescopes from two Indian institutes namely Aryabhatta Research Institute of Observational Sciences (ARIES) and Physical Research Laboratory (PRL) were used for the Intra-night monitoring of the seven RQNLSy1s in the broad-band Johnson-Cousin filter R except for a session with 2.5m PRL telescope when it was taken in the SDSS filter r due to non-availability of broad-band Johnson-Cousin filter R. Broad-band Johnson-Cousin filter R and SDSS filter r were chosen for observations because CCD detector used has maximum sensitivity in these bands. A total of four telescopes two from ARIES namely 1.04 meter (m) Sampurnanand telescope (ST, Sagar, 1999), 1.30m Devasthal Fast Optical Telescope (DFOT, Sagar et al., 2010) located at Nainital, Uttarakhand, and two from PRL namely, 1.2m telescope (Srivastava et al., 2021) and 2.5m telescope located at Mount Abu Rajasthan, were used in this work. The details of the observational set-ups used for each telescope in observing the sample of 7 RQNLSy1s are listed in Table 2. At least three epochs of observation each 3 hours were devoted for each RQNLSy1s. Since we have used in the present work, the optical telescopes range between 1.04m and 2.5m, therefore, the typical exposure time was set between 300 sec to 1200 sec to reach a suitable signal-to-noise ratio (SNR), depending on the sky condition, moon phase, telescope efficiency, and magnitude (brightness) of the RQNLSy1s.
For preliminary processing of the raw images, at least three bias frames, and also three flat frames were taken during each observing session. Furthermore, the standard tasks available in the IRAF222Image Reduction and Analysis Facility (http://iraf.noao.edu/) software package were followed for making final science images from the raw images. Since the field of each target RQNLSy1s was not clustered, therefore, aperture photometry (Stetson, 1987, 1992) was used in the current work for extracting the instrumental magnitudes of RQNLSy1s and the comparison stars registered in the CCD frames, using DAOPHOT II algorithm333Dominion Astrophysical Observatory Photometry (http://www.astro.wisc.edu/sirtf/daophot2.pdf). As emphasized in Ojha et al. (2021) size of the chosen aperture is an important parameter while estimating the instrumental magnitude and the corresponding SNR of the individual photometric data points registered on the CCD frames. Additionally, caution about the point spread function (PSF) variation becomes very important when dealing with intra-night variability of nearby ( 0.4) AGNs. Because in such a situation a significant contribution to the total flux can come from the underlying host galaxy that can mimic the INOV in the standard analysis of the differential light curves (DLCs) due to the significant relative contributions of the (point-like) AGN and the host galaxy to the aperture photometry with the variation of PSF during the session (Cellone et al., 2000). Therefore, the procedure of data reduction, PSF estimation for aperture photometry, selection of aperture, and caution for PSF variations (see Sect. 4) were followed from Ojha et al. (2021). Since except for one RQNLSy1 J102906.69555625.2 (), all the RQNLSy1s in the present sample are at lower redshift ( 0.4), therefore proper caution has been taken about its PSF variation during the night before commenting about its variability (see Sect. 3).
Furthermore, DLCs of target RQNLSy1 for each intra-night session were derived relative to a pair of non-varying (steady) comparison stars (see online Figs. A1), additionally, the PSF variation during each intra-night session is plotted in the bottom panel of DLCs.
3 Methodology for analysis
3.1 Statistical tests
To examine the presence of intra-night variability in the present sample of 7 RQNLSy1s, we have applied two different flavors of F-test, (i) standard -test (-test) and (ii) the power-enhanced -test (-test), by following the basic requirement and procedure of these two tests as described in (Goyal et al., 2012; de Diego, 2014). Several starstar DLCs were generated for each session with the instrumental magnitudes extracted from the aperture photometry, and out of several pairs of starstar DLCs those two stars were chosen as comparison stars for which no-variability resulted based upon -test. Out of the chosen two comparison stars, the one with the closest match (1, a requirement of -test) in magnitude to the target RQNLSy1 is chosen as a reference star, and other as comparison star. Furthermore, two versions of -test are applied to the DLCs of target RQNLSy1 relative to the reference star and comparison star (basic parameters of these two stars are tabulated in the online Table A1). The -values for the two RQNLSy1 DLCs and starstar DLC of an intra-night session can be written as (e.g. Goyal et al., 2012):
where , , and are the variances with , , and being the mean square (formal) rms errors of the i data points in the DLCs of target RQNLSy1, and is the number of observations. A computed value of 0.05 based upon the data of 262 intra-night monitoring sessions of AGNs by Goyal et al. (2013a) is used here for the correct use of rms errors on the photometric data points.
In Column 6 of online Table A2, we compare the computed -values, resulting from Eq. 1 for a session, with its estimated critical value (=) for the same session, here is the level of significance for the test. The values in the current work are set by us to be 0.05 and 0.01, corresponding to 95 percent and 99 percent confidence levels for INOV detection. The null hypothesis (i.e., non-detection of INOV) is discarded at the level of significance if the computed value of exceeds its at the corresponding confidence level. Thus a RQNLSy1 is assigned as a variable (V) if the computed value of is found to be greater than its ; probable variable (PV) if the same is found to be greater than but less or equal to , and non-variable if is found to be less than or equal to . Summary of computed values of and correspondingly inferred status of INOV detection for all the 28 sessions are tabulated in columns 6 and 7 of online Table A2.
The second flavor of -test (the -test) can be written following de Diego (2014) as below
here is the variance of the ‘target RQNLSy1-reference star’ DLC, while is the combined variance of ‘comparison star-reference star’ DLC having data points (number of observations) and comparison stars, computed using scaled square deviation as
where, ’s is the ‘j comparison star-reference star’ differential instrumental magnitudes value and represent the corresponding average value of the DLC for its data points. The scaling factor is taken here as described in Ojha et al. (2021).
Columns 10 and 11 of online Table A2 represent the -test values and corresponding inferred INOV status for the entire session, following the criteria as set for -test (see above).
In addition to variability resulting from two versions of the F-test, proper caution has also been taken about its PSF variation during the night before finalizing its variability status because except for one RQNLSy1 J102906.69555625.2 (), all the RQNLSy1s in the present sample are at lower redshift ( 0.4). Thus for a genuine INOV detection from the current sample, we have first carefully inspected the seeing variations of all the variable (including probable variable cases) intra-night sessions, resulting from -test. An RQNLSy1 is designated as V if either the FWHM of the session was fairly steady during the time of RQNLSy1’s flux variations or gradients in the FWHM of the session are anti-correlating with the systematic variations of differential magnitude of target RQNLSy1 and chosen comparison stars (see Cellone et al., 2000).
3.2 INOV duty cycle estimation
Adopting the definition given by Romero et al. (1999) (see, also Stalin et al., 2004) for the DC of intra-night variability, we have computed it for the current sample of RQNLSy1s with the following expression
where z is the observed duration of monitoring session obtained after redshift correction for the target. For session, is considered to be 1 in Eq. 4 only when INOV is detected, otherwise = 0. The computed INOV duty cycles for the current sample of RQNLSy1s are listed in Table 3, based on two statistical tests .
To compute the peak-to-peak amplitude of INOV () detected in a given DLC, we followed the definition given by Heidt & Wagner (1996)
with = minimum (maximum) values in the DLC of target NLSy1 relative to steady comparison stars and , where, is the mean square (formal) rms errors of individual data points. The mean value of () for different sets (e.g., see Table 3) of RQNLSy1 galaxies is computed by taking the average of the computed values for the DLCs belonging to the “V” category. In Table 3, we have also summarized the DC and values based on the two statistical tests for the different sets of RLNLSy1s, accessed from Ojha et al. (2021, 2022).
Duty cycle and amplitude of INOV () for the individual RQNLSy1 | ||||
-test | -test | |||
RQNLSy1s | DC | DC | ||
J102906.69555625.2 [3] | 31.3 (64.3) | 17.6 (19.5) | 64.3 (64.3) | 19.5 (19.5) |
J122844.81501751.3 [5] | 22.6 (22.6) | 32.7 (32.7) | 38.8 (38.8) | 26.3 (26.3) |
J123220.11495721.8 [5] | 21.0 (21.0) | 29.0 (29.0) | 21.0 (42.2) | 29.0 (21.7) |
J150916.17613716.6 [3] | 00.0 (00.0) | 00.0 (00.0) | 00.0 (00.0) | 00.0 (00.0) |
J151020.05554722.0 [4] | 00.0 (00.0) | 00.0 (00.0) | 25.2 (25.2) | 10.1 (10.1) |
J152205.41393441.3 [4] | 20.1 (20.1) | 09.6 (09.6) | 20.1 (40.7) | 09.6 (07.1) |
J164100.10345452.7 [4] | 00.0 (00.0) | 00.0 (00.0) | 29.6 (00.0) | 11.4 (11.4) |
Duty cycle and amplitude of INOV () for the current and control sample of NLSy1s | ||||
-test | -test | |||
No. of RQNLSy1s | DC | DC | ||
7 RQNLSy1s [28] | 14.3 (17.5) | 19.2 (22.0) | 27.7 (35.0) | 18.9 (17.1) |
15-RLNLSy1s [36] | 30.4 (30.4) | 14.1 (14.1) | 40.5 (47.5) | 13.9 (13.1) |
8 J--RLNLSy1s [23] | 25.9 (25.9) | 08.6 (08.6) | 37.5 (54.7) | 08.1 (07.7) |
The mean value for all the DLCs belonging to the type ‘V’. The number of sessions | ||||
used is tabulated inside the bracket ‘[]’. Values inside parentheses have resulted | ||||
when ‘PV’ cases are considered to ‘V’. | ||||
DC and of 15 radio-loud -ray detected NLSy1s (-RLNLSy1s) are estimated | ||||
using their 36 intra-night sessions from Ojha et al. (2021). | ||||
DC and of eight radio-loud jetted with -ray detected NLSy1s (J--RLNLSy1s) | ||||
are estimated using their 23 intra-night sessions from Ojha et al. (2022). |
4 Results and discussion
The current study presents the first attempt to systematically characterize the INOV for a sample of seven RQNLSy1s that had shown recurring flaring at 37 GHz when observed at MRO (Lähteenmäki et al., 2018). These seven NLSy1s are either radio-quiet or radio-silent (never detected in radio at any frequency). We monitored the current sample (see online Table A2) of seven RQNLSy1s in a total of 28 intra-night sessions, each 3 h (see online Figs. A1). It may be emphasized that an AGN may not show variability on every night it was observed, therefore to improve INOV statistics, we have devoted at least three intra-night sessions, each 3 h for each RQNLSy1s. We applied two versions of the test i.e. -test and -test on the derived 28 intra-night DLCs to confirm the presence/absence of INOV in an intra-night session. Out of 28 intra-night sessions significant INOV with 10% was detected from the DLCs of four RQNLSy1s namely J102906.69555625.2, J122844.81501751.3, J123220.11495721.8, and J152205.41393441.3 (see also first three online figures of Figs. A1) with -test. The DLCs of another session for an RQNLSy1 J102906.69555625.2 showed probable variable (PV) case with the same test even though with 21%. The PV case would have stemmed even though 21%, in this case, may be due to comparatively more noise in its DLCs (see column 12 of online Table A2). Furthermore, eight variable cases with 10% from the DLCs of six RQNLSy1s resulted with -test. Another two sessions with 4% were placed under PV category using the same test. Thus, using statistical tests (see online table Table A2) we find that all RQNLSy1s exhibit INOV except one RQNLSy1, J150916.17613716.6 that did not show INOV with any of the statistical tests for all intra-night sessions. It may be recalled here that out of seven RQNLSy1s observed with JVLA, RQNLSy1s J150916.17613716.6 is the only one with absolutely no detection in the JVLA, while the others are typically showing at least one data point at some frequencies (see Berton et al., 2020b; Järvelä et al., 2023). Therefore, the non-detection of INOV in this source may be due n to its quiescent phase in the optical band too.
From Table 3, a DC of 28% ( 35% when two ‘PV’ cases are considered to be ‘V’) with (the mean value of for all the DLCs belonging to the type ‘V’) of 19% resulted from -test for the current sample. However, DC of 14% ( 18% when a PV case is also considered to be ‘V’) with of 19% are estimated based on the more conservative -test. Considering the average DC () of conservative -test and -test which is 21% ( 26% when ‘PV’ cases are considered to be ‘V), found to be comparable to DC of 25% 30%, exhibited by -RLNLSy1s, that display blazar-like INOV (Ojha et al., 2021). The strong INOV level of 10% in any variable cases resulting from the current sample appear striking because even powerful radio-quiet quasars empowered by SMBH never displayed 4% (e.g., see Gopal-Krishna & Wiita, 2018, and references therein). From Table 3, it also appears that duty cycles of individual variable sources are always 20% with 10%, consistent with the DC estimated for the whole sample. Here, it may be recalled that in the extensive INOV study of six prominent classes of luminous AGNs covered in 262 monitoring sessions by Goyal et al. (2013b) where a very similar telescopes and analysis procedure were used, only blazars class of AGN displayed a DC above 10% for 3%. Thus, we conclude that the DC resulting from the present sample appears to be blazar-like.
Considering the median black hole mass log() = 6.84 that implied to harbor inevitably less powerful jets for the current sample (Heinz & Sunyaev, 2003; Foschini, 2014). The resulting of 21% from the current sample appears striking and it suggests the origin of their INOV from the relativistic plasma jets. However, it may be recalled here that the median black hole mass of the current sample is an order of less massive than the median black hole mass of log(M) = 7.72 of jetted-RLNLSy1s (see last column of Table 5 of Ojha et al., 2022). Thus, despite having low SMBH our sample of NLSy1s is capable of launching relativistic jets. This is in agreement with the recent results of blazar-like INOV displayed by a sample of 12 low-mass (median M) radio-quiet AGNs (Gopal-Krishna et al., 2023), which strengthened the case for the ability to launch relativistic jets from apparently radio-quiet AGNs.
The detection of recurring flaring at 37 GHz at MRO (Lähteenmäki et al., 2018) and strong INOV levels found here from the current sample of seven RQNLSy1s hints at the presence of relativistic jets in the current sample of RQNLSy1s, however, non-detection of jet activity in JVLA observations at 1.6, 5.2, and 9.0 GHz frequencies appears contrasting. Here, it may be recalled that powerful relativistic jets that are capable of propagating outside the host galaxy are found in approximately 10% of AGNs (Padovani, 2017). Therefore, non-detection of jet activity in JVLA observations from the current sample may be due to their low integrated luminosity which is indeed low 1.0 erg s (see Table 2 of Berton et al., 2020b). Such scenario straightens with integrated luminosity 1.5 erg s found in jetted sources (see Berton et al., 2018). In addition to the low power of jets, non-detection of jet activity in JVLA observations might be due to absorbed jets because of a more tied connection of optical emission to nuclear jet emission at millimeter wavelengths as compared to its emission at lower radio frequencies (see Gopal-Krishna et al., 2023), which is largely attenuated due to a high opacity around the nuclear jet, as interpreted from Very Long Baseline Interferometry studies (Gopal-Krishna & Steppe, 1991; Boccardi et al., 2017).
The detection of multiple flaring at 37 GHz and INOV in the current work while non-detection of jet activity in JVLA observations at 1.6, 5.2, and 9.0 GHz frequencies may be due to the flaring and quiescent states of RQNLSy1s when 37 GHZ, our observations, and JVLA observations were taken place, respectively.
Another possibility of non-detection of jet activity in JVLA observations as also emphasised in Berton et al. (2020b) from the current sample could be due to their high-frequency peakers nature which usually happens with extremely young objects like NLSy1s which are expected to be in an early phase of their evolution (see Komossa, 2018; Paliya, 2019). It may be emphasised here that NLSy1s are young and typically characterized by high Eddington ratios (Boroson & Green, 1992; Ojha et al., 2020), therefore expected to be associated with a dense circumnuclear environment around non-flattened broad-line region (Heckman & Best, 2014; Vietri et al., 2018; Berton et al., 2020a) that might also hinder relativistic jets propagation through its interaction with the clouds (vanBreugel et al., 1984).
In addition to non-detection of jet activity in the JVLA observations at 1.6, 5.2, and 9.0 GHz, recent observations with JVLA at higher frequencies 10, 15, 22, 33, and, 45 GHz showed no signs of jets from the current sample rather than resulting in either steep spectrum or no detection at all from most of the sources except for a source J122844.81501751.2 that showed the flat spectrum (Järvelä et al., 2023). However, blazar-like INOV found here from the current sample of RQNLSy1s could be due to the peculiar geometry of jet in these sources that causes changes in viewing angle toward the observer’s line of sight thus changes in Doppler factor (Raiteri et al., 2017).
Finally, one last possibility of non-detection of jet activity in JVLA observations might be magnetic reconnection in the magnetosphere of black hole (Lazarian & Vishniac, 1999; deGouveia Dal Pino et al., 2010; Kadowaki et al., 2015; Ripperda et al., 2022; Kimura et al., 2022) which does not require the presence of a permanent relativistic jet and can account timescales of variability ranging from minutes to days. In brief, this model invokes the interactions of magnetic field lines emerging from the accretion disk with the magnetosphere anchored into the central black hole horizon (Blandford & Znajek, 1977). With the enhancement of the accretion rate, magnetic fluxes from the accretion disc and those anchored into the black hole horizon are pushed together in the inner disk region and reconnected under finite magnetic resistivity (see Figure 1 of Kadowaki et al., 2015). This reconnection becomes very efficient and fast under turbulence instability and releases a huge amount of magnetic power. A part of released magnetic power accelerates particles to relativistic velocities and thus is attributed to radio emissions and flares. Therefore, recurring flaring at 37 GHz observed at MRO might be due to fast magnetic reconnection driven by turbulence. Thus, it implies that jets may not be present in our sample as evidenced from their recent radio studies by JVLA and VLBA (see Berton et al., 2020a; Järvelä et al., 2023), but the INOV seen in the current study may be due to magnetic reconnection in the black hole magnetosphere. Furthermore, for the low luminosity AGNs, deGouveia Dal Pino et al. (2010) and Kadowaki et al. (2015) showed that the turbulence-induced fast magnetic reconnection events in an efficiently accreting black hole with a mass of M can release to erg s power which is sufficient to explain the detection of INOV and flares. The jetted AGN such as gamma-ray NLSy1s and blazars can also possess magnetic reconnection and instabilities in their accretion disks, however, jet-dominated non-thermal emission completely overwhelms the thermal emission from the accretion disc in them (see Mangalam & Wiita, 1993; Ulrich et al., 1997; Blandford et al., 2019).
Additionally, magnetic reconnection events causing acceleration of plasma to relativistic velocities can also potentially heat the corona of the accretion disk resulting in the enhancement of thermal X-ray emission. Furthermore, emissions in high energy gamma-rays might also be possible through interactions of these accelerated relativistic electrons with photon density in the surrounds of the black hole via SSC and/or IC interactions (Kadowaki et al., 2015). Therefore, coordinated radio, X-ray, and gamma-ray observations can be useful in confirming or ruling out the possibility of a magnetic reconnection mechanism (deGouveia Dal Pino et al., 2010; Kimura et al., 2022).
5 Conclusions
In the current study, we present the first attempt to systematically characterize the INOV for a sample of seven RQNLSy1s that had shown recurring flaring at 37 GHz when observed with MRO (Lähteenmäki et al., 2017, 2018), however, no signs of jet were detected from them when these RQNLSy1s were observed in radio bands with JVLA at low frequencies. Recurring flaring at 37 GHz from them strongly suggests the presence of jets in them. However, the non-detection of jet activity in JVLA and VLBA observations appears puzzling. We have addressed this issue by taking advantage of INOV which is being used to infer the presence of relativistic jets in AGNs based on their blazar-like duty cycle and the amplitude of INOV. Therefore, we monitored the current sample in a total of 28 intra-night sessions, each 3 h. The resulting level of INOV from this sample found to be similar to the INOV level of -RLNLSy1s, which displays blazar-like INOV. Thus detection of recurring flaring at 37 GHz with MRO and strong INOV level found here from the current sample of seven RQNLSy1s hints at the presence of relativistic jets in the present set of RQNLSy1s. Furthermore, the resulting strong level of INOV from the current sample of low-mass RQNLSy1s and almost a similar INOV level observed recently by Gopal-Krishna et al. (2023) from 12 low-mass active galactic nuclei, suggest that even low-mass radio-quiet and/or radio-silent AGNs can launch relativistic jets. Furthermore, inferred jet activity from the current sample along with the presence of jet activity in 12 low-mass AGNs (Gopal-Krishna et al., 2023) would be useful in understanding relativistic jet mechanism in lower black hole mass (M) AGNs.
The resulting variability nature of these sources in radio and optical wavebands is difficult to explain with the usual variability mechanisms of AGNs. Therefore, studying such types of sources with bigger sample sizes is very important to unveil the nature of these sources as they might represent a new type of AGN variability.
Acknowledgements
We are thankful to the anonymous referee for providing comments and suggestions, which helped us to improve the manuscript considerably. We thank Dr. Mudit Srivastava and Dr. Vipin Kumar for their help in carrying out observations with MFOSC-P on the 1.2m telescope and also acknowledge the help of Dr. Vishal Joshi for using FOC on the 2.5m telescope. We also thank the concerned ARIES and PRL staff for assistance during the observations obtained at the 1.04m Sampurnanand Telescope (ST), 1.30m Devasthal Fast Optical Telescope (DFOT) of ARIES and 1.20m, 2.50m Mt. Abu telescopes of PRL, which are the national facilities run and managed by ARIES and PRL, respectively.
Data availability
The data from the ARIES telescopes and the Mount Abu observatory Mt. Abu telescopes of PRL used in this paper will be shared on reasonable request to the corresponding author.
Appendix A Appendix
RQNLSy1s and | Date(s) of monitoring | RA.(J2000) | DEC.(J2000) | g | r | g-r |
the comparison stars | (hh mm ss) | (dd mm ss) | (mag) | (mag) | (mag) | |
(1) | (2) | (3) | (4) | (5) | (6) | (7) |
J102906.69555625.2 | 2020 Dec. 19; 2021 Apr. 12; 2022 Dec. 27 | 10:29:06.69 | 55:56:25.22 | 19.08 | 19.07 | 0.01 |
S1 | 2020 Dec. 19; 2022 Dec. 27 | 10:28:41.50 | 55:57:21.89 | 19.69 | 18.34 | 1.35 |
S2 | 2020 Dec. 19; 2022 Dec. 27 | 10:28:21.48 | 55:58:34.46 | 19.38 | 18.18 | 1.20 |
S3 | 2021 Apr. 12 | 10:29:57.00 | 55:54:27.70 | 18.59 | 17.67 | 0.92 |
S4 | 2021 Apr. 12 | 10:29:54.51 | 55:55:04.26 | 19.25 | 17.82 | 1.43 |
J122844.81501751.3 | 2020 Dec. 23; 2021 Apr. 10; 2022 Mar. 27, Apr. 09, May 08 | 12:28:44.81 | 50:17:51.27 | 18.44 | 17.88 | 0.56 |
S1 | 2020 Dec. 23 | 12:28:20.67 | 50:22:08.95 | 18.04 | 17.28 | 0.76 |
S2 | 2020 Dec. 23 | 12:28:58.53 | 50:20:11.89 | 16.55 | 15.75 | 0.80 |
S3 | 2021 Apr. 10 | 12:28:34.32 | 50:15:40.12 | 17.10 | 16.30 | 0.80 |
S4 | 2021 Apr. 10, 2022 Apr. 09 | 12:28:48.49 | 50:15:26.72 | 17.46 | 16.60 | 0.86 |
S5 | 2022 Mar. 27, May 08 | 12:28:54.32 | 50:16:22.36 | 18.64 | 17.39 | 1.25 |
S6 | 2022 Mar. 27, May 08 | 12:28:45.45 | 50:14:48.60 | 17.58 | 17.21 | 0.37 |
S7 | 2022 Apr. 09 | 12:28:45.32 | 50:18:40.78 | 18.11 | 17.49 | 0.73 |
J123220.11495721.8 | 2021 Jan. 24, Apr. 11; 2022 Jan 14, 23, 24 | 12:32:20.11 | 49:57:21.79 | 17.69 | 17.54 | 0.62 |
S1 | 2021 Jan. 24, Apr. 11; 2022 Jan 14 | 12:31:59.11 | 49:58:50.29 | 17.65 | 16.24 | 1.41 |
S2 | 2021 Jan. 24 | 12:32:55.82 | 49:52:58.96 | 18.08 | 17.23 | 0.85 |
S3 | 2021 Apr. 11; 2022 Jan. 14, 23, 24 | 12:32:15.53 | 49:56:07.70 | 17.15 | 15.93 | 1.22 |
S4 | 2022 Jan 23, 24 | 12:32:09.86 | 49:56:08.93 | 17.35 | 17.03 | 0.32 |
J150916.17613716.6 | 2021 Apr. 08, 10; 2022 Mar. 27 | 15:09:16.17 | 61:37:16.75 | 18.88 | 18.47 | 0.41 |
S1 | 2021 Apr. 08, 10 | 15:09:07.14 | 61:37:45.68 | 17.41 | 16.96 | 0.45 |
S2 | 2021 Apr. 08, 10 | 15:09:32.71 | 61:37:57.88 | 17.71 | 16.66 | 1.05 |
S3 | 2022 Mar. 27 | 15:09:38.99 | 61:34:36.71 | 18.56 | 17.42 | 1.14 |
S4 | 2022 Mar. 27 | 15:09:19.99 | 61:32:17.34 | 16.80 | 16.43 | 0.37 |
J151020.05554722.0 | 2021 Apr. 07, 11; 2022 Apr. 10, 23 | 15:10:20.05 | 55:47:22.05 | 18.46 | 17.77 | 0.69 |
S1 | 2021 Apr. 07 | 15:09:44.54 | 55:44:07.17 | 18.27 | 17.27 | 1.00 |
S2 | 2021 Apr. 07, 11 | 15:09:59.61 | 55:51:28.95 | 17.38 | 16.98 | 0.40 |
S3 | 2021 Apr. 11; 2022 Apr. 10 | 15:10:02.94 | 55:49:18.50 | 16.95 | 15.91 | 1.04 |
S4 | 2021 Apr. 11; 2022 Apr. 10, 23 | 15:10:17.66 | 55:47:41.71 | 17.65 | 16.72 | 0.93 |
S5 | 2022 Apr. 23 | 15:10:17.29 | 55:50:47.03 | 15.90 | 15.49 | 0.41 |
J152205.41393441.3 | 2021 Jun. 01, 02, 03; 2022 Feb. 21 | 15 22 05.41 | 39 34 41.30 | 14.90 | 13.10 | 1.80 |
S1 | 2021 Jun. 01, 02 | 15:22:15.47 | 39:30:41.30 | 15.50 | 15.49 | 0.01 |
S2 | 2021 Jun. 01, 02 | 15:22:15.07 | 39:36:13.70 | 16.90 | 15.50 | 1.40 |
S3 | 2021 Jun. 03 | 15:21:22.87 | 39:35:07.20 | 15.80 | 15.20 | 0.60 |
S4 | 2021 Jun. 03 | 15:21:18.16 | 39:35:49.60 | 15.50 | 15.20 | 0.30 |
S5 | 2022 Feb. 21 | 15:22:02.09 | 39:31:31.10 | 15.70 | 15.69 | 0.01 |
S6 | 2022 Feb. 21 | 15:21:53.28 | 39:32:35.40 | 14.30 | 14.20 | 0.10 |
J164100.10345452.7 | 2021 June, 01, 06; 2022 Apr. 23, May 07 | 16:41:00.10 | 34:54:52.68 | 17.94 | 16.96 | 0.98 |
S1 | 2021 June, 01 | 16:41:07.53 | 34:49:34.97 | 17.26 | 16.09 | 1.17 |
S2 | 2021 June, 01, 06 | 16:40:39.77 | 34:57:31.21 | 16.96 | 15.98 | 0.98 |
S3 | 2021 June, 06 | 16:40:37.99 | 35:00:33.55 | 16.21 | 15.64 | 0.57 |
S4 | 2022 Apr. 23, May 07 | 16:41:10.04 | 34:51:09.22 | 16.47 | 16.10 | 0.37 |
S5 | 2022 Apr. 23, May 07 | 16:41:00.67 | 34:53:31.39 | 17.37 | 15.98 | 1.39 |
RQNLSy1s | Date(s) | T | N | Median | -test | INOV | -test | Variability | -test | INOV | ||
(SDSS name) | yyyy.mm.dd | (hrs) | FWHM | , | status | status of | status | (AGN-s) | (%) | |||
(arcsec) | 99% | 99% | (s1s2) | 99% | ||||||||
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) | (12) | (13) |
J102906.69555625.2 | 2020.12.19 | 4.80 | 17 | 2.56 | 03.40, 03.41 | V, V | 00.86 | NV | 03.64 | V | 0.03 | 17.62 |
2021.04.12 | 4.55 | 14 | 2.22 | 03.06, 02.62 | PV, PV | 00.77 | NV | 03.96 | V | 0.04 | 21.33 | |
2022.12.27 | 4.21 | 22 | 2.27 | 00.16, 00.14 | NV, NV | 00.29 | NV | 00.53 | NV | 0.05 | – | |
J122844.81501751.3 | 2020.12.23 | 3.49 | 22 | 3.53 | 00.71, 01.09 | NV, NV | 00.81 | NV | 00.88 | NV | 0.02 | – |
2021.04.10 | 3.10 | 36 | 2.30 | 00.61, 00.53 | NV, NV | 00.64 | NV | 00.96 | NV | 0.04 | – | |
2022.03.27 | 4.49 | 59 | 1.79 | 00.44, 00.34 | NV, NV | 00.35 | NV | 01.25 | NV | 0.04 | – | |
2022.04.09 | 3.26 | 32 | 1.87 | 02.38, 02.40 | V, V | 00.67 | NV | 03.21 | V | 0.05 | 32.67 | |
2022.05.08 | 4.53 | 17 | 1.90 | 00.59, 00.55 | NV, NV | 00.14 | NV | 04.37 | V | 0.07 | 19.99 | |
J123220.11495721.8 | 2021.01.24 | 3.25 | 19 | 1.86 | 00.71, 00.50 | NV, NV | 00.41 | NV | 01.73 | NV | 0.01 | – |
2021.04.11 | 3.69 | 43 | 2.33 | 00.35, 00.34 | NV, NV | 00.57 | NV | 00.61 | NV | 0.02 | – | |
2022.01.14 | 3.04 | 15 | 1.92 | 02.23, 01.90 | NV, NV | 02.10 | NV | 01.06 | NV | 0.02 | - - | |
2022.01.23 | 3.03 | 17 | 2.44 | 05.44, 05.54 | V, V | 00.87 | NV | 06.27 | V | 0.03 | 28.96 | |
2022.01.24 | 3.00 | 30 | 2.18 | 01.50, 01.62 | NV, NV | 00.68 | NV | 02.20 | PV | 0.03 | (14.36) | |
J150916.17613716.6 | 2021.04.08 | 3.20 | 13 | 2.71 | 01.83, 01.92 | NV, NV | 01.84 | NV | 00.99 | NV | 0.03 | – |
2021.04.10 | 4.02 | 16 | 2.18 | 00.93, 00.95 | NV, NV | 00.52 | NV | 01.79 | NV | 0.03 | – | |
2022.03.27 | 3.11 | 12 | 2.10 | 01.12, 01.50 | NV, NV | 01.11 | NV | 01.01 | NV | 0.04 | – | |
J151020.05554722.0 | 2021.04.07 | 3.14 | 18 | 1.86 | 00.71, 00.84 | NV, NV | 00.56 | NV | 01.28 | NV | 0.02 | – |
2021.04.11 | 3.45 | 14 | 2.44 | 02.25, 02.35 | NV, NV | 00.46 | NV | 04.85 | V | 0.02 | 10.13 | |
2022.04.10 | 4.16 | 11 | 2.40 | 00.44, 00.61 | NV, NV | 00.99 | NV | 00.44 | NV | 0.03 | – | |
2022.04.23 | 3.30 | 10 | 2.41 | 00.95, 00.33 | NV, NV | 01.22 | NV | 00.77 | NV | 0.04 | – | |
J152205.41393441.3 | 2021.06.01 | 3.10 | 36 | 2.68 | 00.93, 01.10 | NV, NV | 00.68 | NV | 01.38 | NV | 0.01 | – |
2021.06.02 | 3.81 | 35 | 3.06 | 03.65, 04.78 | V, V | 01.14 | NV | 03.21 | V | 0.01 | 09.60 | |
2021.06.03 | 3.70 | 41 | 2.70 | 01.52, 01.49 | NV, NV | 00.81 | NV | 01.89 | PV | 0.01 | (04.57) | |
2022.02.21 | 2.21 | 23 | 2.48 | 01.16, 02.34 | NV, NV | 02.02 | NV | 00.51 | NV | 0.01 | – | |
J164100.10345452.7 | 2021.06.01 | 3.00 | 29 | 3.16 | 00.87, 00.74 | NV, NV | 00.49 | NV | 01.78 | NV | 0.02 | – |
2021.06.06 | 6.95 | 67 | 2.82 | 00.77, 00.66 | NV, NV | 00.63 | NV | 01.22 | NV | 0.02 | – | |
2022.04.23 | 3.23 | 13 | 2.30 | 01.64, 02.09 | NV, NV | 01.56 | NV | 01.05 | NV | 0.01 | – | |
2022.05.07 | 3.02 | 10 | 2.54 | 03.14, 02.92 | NV, NV | 00.35 | NV | 08.94 | V | 0.02 | 11.45 | |
Date(s) of the monitoring session(s). Duration of the monitoring session in the observed frame. Number of data points in the DLCs of the monitoring session. | ||||||||||||
Median seeing (FWHM in arcsec) for the session. Variability status inferred from F and F tests, with V = variable, i.e. confidence level 99%; | ||||||||||||
PV = probable variable, i.e. % confidence level; NV = non-variable, i.e. confidence level 95%. | ||||||||||||
Mean amplitude of variability in the two DLCs of the target RQNLSy1 (i.e., relative to the two chosen comparison stars). |